# A Note on Output Length of One-Way State Generators

### v1

@@ 1.2 Discussion
Our results suggest that the upper and lower bounds for output length for ($\epsilon$-)OWSGs behave quite differently from those for OWFs.
  • It is easy to see that weak OWFs with $O(\log\lambda)$-bit outputs do not exist. On the other hand, assuming the existence of exponentially secure OWFs with linear expansion, we can show that there exist (strong) OWFs with $m$-bit outputs for any $m=\omega(\log\lambda)$ by a simple padding argument. This means that $O(\log\lambda)$ is the best lower bound for both OWFs and weak OWFs. On the other hand, our results suggest that the lower bounds for OWSGs and weak OWSGs are different. In particular, we show that there do not exist OWSGs with $O(\log\lambda)$-qubit outputs (Theorem 1.1), but there exist weak OWSGs with slightly super-constant-qubit outputs (Theorem 1.4), which circumvents the barrier for OWSGs.
  • For constructing OWFs with the optimal output length, i.e., those with slightly super-logarithmic-bit outputs, we need to assume exponential hardness of $\mathbf{NP}$ due to an obvious reason. On the other hand, we can construct OWSGs with slightly super-logarithmic-bit outputs only assuming polynomial hardness of OWFs [BS20, ${}^{+}$23]. Since we show that their construction is optimal in terms of output length (Theorem 1.1), this means that polynomial hardness is sufficient for achieving the optimal output length for OWSGs.
Our upper and lower bounds of the output length of OWSGs also imply those for several other primitives.
@@
Though we give matching upper and lower bounds for the output length of OWSGs and weak OWSGs, there is still a gap for $\Theta(1)$-OWSGs. In particular, our lower bound is $(\log\log\lambda)/2+O(1)$ (Theorem 1.3) whereas upper bound is $O(\log\log\lambda)$ (Theorem 1.2). It is an interesting open problem to study the parameter regime of constants $\epsilon$ and $c$ for which $\epsilon$-OWSGs with $c\log\log\lambda$-qubit output do and do not exist.
In this paper, we focus on $\epsilon$-OWSGs for $\epsilon\in\{{\mathsf{negl}}(\lambda),\Theta(1),1-1/{\mathrm{poly}}(\lambda)\}$, but another interesting parameter regime is $\epsilon=1/{\mathrm{poly}}(\lambda)$ for some fixed polynomial ${\mathrm{poly}}$. As for a lower bound, by using Welch bound similarly to the proof of Theorem 1.1, we would be able to show that for any polynomial ${\mathrm{poly}}$, there is a constant $c$ such that there do not exist $1/{\mathrm{poly}}(\lambda)$-OWSGs with $c\log\lambda$-qubit outputs. This still leaves a possibility that for any polynomial ${\mathrm{poly}}$, there exist $1/{\mathrm{poly}}(\lambda)$-OWSGs with $O(\log\lambda)$-qubit outputs, which circumvents the barrier for standard OWSGs. On the other hand, if we apply our construction based on quantum fingerprinting similarly to Theorem 1.2, we obtain $1/{\mathrm{poly}}(\lambda)$-OWSGs with $O((\log\lambda)(\log\log\lambda))$-qubit outputs assuming subexponentially secure OWFs. However, this is meaningless since existing works [BS20, ${}^{+}$23] already imply them based on a weaker assumption of standard OWFs. In the quantum random oracle model (QROM) [${}^{+}$11], there is quantum fingerprinting with better output length (see Remark 2.9), which leads to $1/{\mathrm{poly}}(\lambda)$-OWSGs with $O(\log\lambda)$-qubit outputs. We leave it open to construct $1/{\mathrm{poly}}(\lambda)$-OWSGs with similar parameters in the plain model (i.e., without relying on random oracles).
Another interesting question is to construct quantum OWFs using short output OWSGs. Here, a quantum OWF is a quantum (pseudo)deterministic polynomial-time algorithm whose inputs and outputs are classical bit strings and the standard one-wayness is satisfied. The construction of quantum PRGs from short output PRSGs in [ALY23] would not be applied to the case of OWSGs, because both the proofs for (pseudodeterministic) correctness and security exploit (pseudo)randomness of output states of PRSGs. Answering this question provides a better understanding of quantum cryptographic primitives with short output lengths, and simplifies the proof of Theorems 1.3 and 1.5.
@@ 1.3 Organization
We prepare definitions and technical tools in Section 2. We prove our results on lower bounds for output length of ($\epsilon$-)OWSGs (Theorems 1.1, 1.5 and 1.3) in Section 3. We prove our results on upper bounds for output length of ($\epsilon$-)OWSGs (Theorems 1.2 and 1.4) in Section 4.
@@ 2 Preliminaries
A function $f:\{0,1\}^{*}\to\{0,1\}^{*}$ is a (quantum-secure) OWF if it is computable in classical deterministic polynomial-time, and for any QPT adversary $\mathcal{A}$, there exists a negligible function ${\mathsf{negl}}$ such that for any $\lambda$,
@@
Definition 2.3 (One-way state generators (OWSGs)).
@@
We use the following version of quantum state tomography given in [ALY23].
@@
Let $\delta=\delta(\lambda)\in(0,1]$ and $d=d(\lambda)\in\mathbb{N}$. There exists a tomography procedure $\mathsf{Tomography}$ that takes $t(\lambda):=144\lambda d^{4}/\delta^{2}$ copies of a $d$-dimensional mixed quantum state $\rho$, and outputs a matrix $M\in\mathbb{C}^{d\times d}$ such that
@@
Remark 2.9 ().
The work [BCWdW01, Section 3] also shows an alternative non-constructive method for quantum fingerprinting with better output length. With some rephrasing, they show that we can take $m=\log\ell+O(\log\eta^{-1})$ in such a way that if we define |$\ket{h_{x}}:=2^{-m/2}\sum_{z\in\{0,1\}^{m}}(-1)^{H(x||z)}\ket{z}$ for a random function $H:\{0,1\}^{\ell+m}\rightarrow\{0,1\}$, then for $(1-2^{-\Omega(\ell)})$-fraction of $H$, we have $|\langle h_{x}|h_{x^{\prime}}\rangle|\leq\eta$ for all pairs of distinct $\ell$-bit strings $x\neq x^{\prime}$. Then they argue that there exists a fixed $H$ which gives a quantum fingerprinting with the desired parameter by a non-constructive argument. We observe that the above construction can be directly used as quantum fingerprinting in the quantum random oracle model (QROM) [${}^{+}$11]. This means that we have a quantum fingerprinting with output length $\log\ell+O(\log\eta^{-1})$ in the QROM, which is better than the one given in Lemma 2.8. We do not know how to achieve the same bound in the plain model (i.e., without relying on random oracles). The improvement from $O(\log(\ell)\log(\eta^{-1}))$ to $\log\ell+O(\log\eta^{-1})$ does not lead to any asymptotical improvement for output length of $\epsilon$-OWSGs in the parameter regimes which we mainly consider in this paper, but this gives an asymptotical improvement for $\epsilon=1/{\mathrm{poly}}(\lambda)$ (see the second open problem in Section 1.2).
3 Lower Bounds
In this section, we prove lower bounds for output length of ($\epsilon$-)OWSGs (Theorems 1.1, 1.5 and 1.3).
@@ 3 Lower Bounds
where Equation 16 follows from the correctness of $(\mathsf{KeyGen},\mathsf{StateGen},\mathsf{Ver})$ and Equation 7. Note that we have $\sqrt{1-x}<1-x/4$ for all $0\leq x\leq\frac{1}{2}$ and $\sqrt{1-x}\leq\sqrt{\frac{1}{2}}$ for all $\frac{1}{2}\leq x\leq 1$. Therefore we have
$\Pr[\mathcal{A}~{}\text{wins}]\geq\min\left\{2^{-m-2},1-\sqrt{\frac{1}{2}}\right\}-{\mathsf{negl}}(\lambda).$ (17)
@@
Because of Lemmata 2.6 and 2.7, the time complexity of this algorithm is bounded by a polynomial in $1/\gamma,d,\lambda,T$ and $|\mathsf{Net}(d,\gamma)$. By the condition, $|\mathsf{Net}(d,\gamma)|\leq(C/\gamma)^{d^{2}}=(6C/\Delta)^{2^{2m}}={\mathrm{poly}}(\lambda)$ and this implies that $1/\Delta={\mathrm{poly}}(\lambda)$ and $d=2^{m}={\mathrm{poly}}(\lambda)$. This proves that $\mathcal{A}$ is a QPT algorithm.
@@
4 Upper Bounds
In this section, we prove upper bounds for output length of ($\epsilon$-)OWSGs (Theorems 1.2 and 1.4).
@@ 4 Upper Bounds
According to [BS20], if there exist OWFs, then there exist PRSGs with input length $\lambda$ and any output length $m$. From [${}^{+}$23, Theorem 3.4], these PRSGs with $m=\omega(\log\lambda)$ are also OWSGs. ∎

### v3

@@ 1.2 Discussion
Our results suggest that the upper and lower bounds for the output length of ($\epsilon$-)OWSGs behave quite differently from those of OWFs.
  • It is easy to see that weak OWFs with $O(\log\lambda)$-bit outputs do not exist. On the other hand, assuming the existence of exponentially secure OWFs with linear expansion, we can show that there exist (strong) OWFs with $m$-bit outputs for any $m=\omega(\log\lambda)$ by a simple padding argument. This means that $O(\log\lambda)$ is the best lower bound for both OWFs and weak OWFs. On the other hand, our results suggest that the lower bounds for OWSGs and weak OWSGs are different. In particular, we show that there do not exist OWSGs with $O(\log\lambda)$-qubit outputs (Theorem 1.1), but there exist weak OWSGs with slightly super-constant-qubit outputs (Theorem 1.7), which circumvents the length barrier for standard OWSGs (Theorem 1.1).
  • For constructing OWFs with the optimal output length, i.e., those with slightly super-logarithmic-bit outputs, we need to assume exponential hardness of $\mathbf{NP}$ due to an obvious reason. On the other hand, we can construct OWSGs with slightly super-logarithmic-bit outputs only assuming polynomial hardness of OWFs [BS20, CGG+23]. Since we show that their construction is optimal in terms of output length (Theorem 1.1), this means that polynomial hardness is sufficient for achieving the optimal output length for OWSGs.
Our upper and lower bounds for the output length of OWSGs also imply those for several other primitives.
@@
Though we give matching upper and lower bounds for the output length of OWSGs and weak OWSGs, there are gaps for $\lambda^{\Theta(1)}$-OWSGs and $\Theta(1)$-OWSGs. For $\lambda^{c}$-OWSGs with $c=\Theta(1)$, our lower bound is $c\log\lambda-2$ (Theorem 1.2) whereas our upper bound is $(c+1)\log\lambda+O(1)$ (Theorem 1.5). For $\Theta(1)$-OWSGs, our lower bound is $(\log\log\lambda)/2+O(1)$ (Theorem 1.3) whereas upper bound is $O(\log\log\lambda)$ (Theorem 1.6). It is an interesting open problem to fill the gap.
Another interesting question is to construct quantum OWFs using short output OWSGs. Here, a quantum OWF is a quantum (pseudo)deterministic polynomial-time algorithm whose inputs and outputs are classical bit strings and the standard one-wayness is satisfied. The construction of quantum PRGs from short output PRSGs in [ALY23] would not be applied to the case of OWSGs, because both the proofs for (pseudodeterministic) correctness and security exploit (pseudo)randomness of output states of PRSGs. Answering this question provides a better understanding of quantum cryptographic primitives with short output lengths, and simplifies the proofs of Theorems 1.3 and 1.4.
@@ 1.3 Organization
We prepare definitions and technical tools in Section 2. We prove our results on lower bounds for the output length of ($\epsilon$-)OWSGs (Theorems 1.1, 1.2, 1.4 and 1.3) in Section 3. We prove our results on upper bounds for the output length of ($\epsilon$-)OWSGs (Theorems 1.6, 1.5 and 1.7) in Section 4. We show our results on lower and upper bounds for the length of EFIs (Theorems 1.8 and 1.9) in Section 5.
@@ 2.1 Basic Cryptographic Primitives
A function $f:\{0,1\}^{*}\to\{0,1\}^{*}$ is a (quantum-secure) OWF if it is computable in classical deterministic polynomial-time, and for any QPT adversary $\mathcal{A}$,
@@
Definition 2.3 (PRGs).
A function $G:\{0,1\}^{*}\to\{0,1\}^{*}$ is a (quantum-secure) PRG if it is computable in classical deterministic polynomial-time, it maps $\lambda$-bit inputs to $\ell(\lambda)$-bit outputs for $\ell(\lambda)\geq\lambda+1$ and for any QPT adversary $\mathcal{A}$,
$\left|\Pr_{x\leftarrow\{0,1\}^{\lambda}}[1\leftarrow\mathcal{A}(1^{\lambda},G(x))]-\Pr_{y\leftarrow\{0,1\}^{\ell}}[1\leftarrow\mathcal{A}(1^{\lambda},y)]\right|\leq{\mathsf{negl}}(\lambda).$ (4)
We say that a PRG is exponentially secure if there exists a constant $10<c<$ 1 such that for any QPT adversary $\mathcal{A}$,
$\left|\Pr_{x\leftarrow\{0,1\}^{\lambda}}[1\leftarrow\mathcal{A}(1^{\lambda},G(x))]-\Pr_{y\leftarrow\{0,1\}^{\ell}}[1\leftarrow\mathcal{A}(1^{\lambda},y)]\right|\leq 2^{-c\lambda}.$ (5)
Definition 2.4 (One-way state generators (OWSGs)  [MY22a]).
@@
Definition 2.5 (Pseudo-random quantum states generators (PRSGs) [JLS18]).
For a function $\epsilon:\mathbb{N}\to[0,1]$, An $\epsilon$-pseudorandom quantum states generator ($\epsilon$-PRSG) is a tuple of QPT algorithms $(\mathsf{KeyGen},\mathsf{StateGen})$ that work as follows:
  • $\mathsf{StateGen}(k)\to\ket{\phi_{k}}:$ It is a QPT algorithm that, on input $k$, outputs an $m$-qubit quantum state |$\ket{\phi_{k}}$.
We require the following pseudorandomness: For any polynomial $t$ and any QPT algorithm $\mathcal{A}$,
$\left|\Pr_{k\leftarrow\mathsf{KeyGen}(1^{\lambda})}\left[1\leftarrow\mathcal{A}\left(1^{\lambda},\ket{\phi_{k}}^{\otimes t}\right)\right]-\Pr_{\ket{\psi}\leftarrow\mu_{m}}\left[1\leftarrow\mathcal{A}\left(1^{\lambda},\ket{\psi}^{\otimes t}\right)\right]\right|\leq\epsilon(\lambda)$
for all sufficiently large $\lambda$, where $\mu_{m}$ is the Haar measure on $m$-qubit quantum states. We say that it is a PRSG if it is a $\lambda^{-c}$-PRSG for all $c\in\mathbb{N}$.
Definition 2.6 (EFIs [BCQ23]).
An EFI is an algorithm $\mathsf{StateGen}(b,1^{\lambda})\to\rho_{b}$ that, on input $b\in\{0,1\}$ and the security parameter $\lambda$, outputs a quantum state $\rho_{b}$ such that all of the following three conditions are satisfied.
  • It is a uniform QPT algorithm.
  • $\rho_{0}$ and $\rho_{1}$ are computationally indistinguishable. In other words, for any QPT adversary $\mathcal{A}$, $|\Pr[1\leftarrow\mathcal{A}(1^{\lambda},\rho_{0})]-\Pr[1\leftarrow\mathcal{A}(1^{\lambda},\rho_{1})]|\leq{\mathsf{negl}}(\lambda).$
  • $\rho_{0}$ and $\rho_{1}$ are statistically distinguishable, i.e., $\|\rho_{0}-\rho_{1}\|_{tr}\geq\frac{1}{{\mathrm{poly}}(\lambda)}.$
We call it an $\ell(\lambda)$-qubit EFI if $\rho_{0}$ and $\rho_{1}$ are $\ell(\lambda)$-qubit states.
Remark 2.7.
In the original definition of EFIs in [BCQ23], computational indistinguishability is required to hold against non-uniform QPT adversaries whereas we only require computational indistinguishability against uniform QPT adversaries. This only makes our result on the lower bound for the output length of EFIs stronger.
2.2 Useful Lemmata
@@
Let $\delta=\delta(\lambda)\in(0,1]$ and $d=d(\lambda)\in\mathbb{N}$. There exists a tomography procedure $\mathsf{Tomography}$ that takes $t(\lambda):=144\lambda d^{4}/\delta^{2}$ copies of a $d$-dimensional mixed quantum state $\rho$, and outputs a Hermitian matrix $M\in\mathbb{C}^{d\times d}$ such that
@@
Remark 2.11.
[ALY23, Corollary 2.16] does not mention that the output matrix is Hermitian. This fact can be checked by following the proof; the starting point is the folklore tomography (see [Low21, Section 1.5.3]) that approximates $\rho=\frac{1}{d}\sum\mathrm{Tr}(P_{i}\rho)P_{i}$ for Pauli matrices $P_{i}$ by $N=\sum\mu_{i}P_{i}$ for $\mu_{i}\approx\frac{1}{d}\mathrm{Tr}(P_{i}\rho)$. The matrix $N$ is Hermitian, and the above lemma, which is a trace-distance version of [AGQY22, Corollary 7.6], only chooses a nice $M$ among many independent $N$’s, so that $M$ is still Hermitian.
@@
Remark 2.14.
The work [BCWdW01, Section 3] also shows an alternative non-constructive method for quantum fingerprinting with better output length. With some rephrasing, they show that we can take $m=\log\ell+O(\log\eta^{-1})$ in such a way that if we define |$\ket{h_{x}}:=2^{-m/2}\sum_{z\in\{0,1\}^{m}}(-1)^{H(x||z)}\ket{z}$ for a random function $H:\{0,1\}^{\ell+m}\rightarrow\{0,1\}$, then for $(1-2^{-\Omega(\ell)})$-fraction of $H$, we have $|\langle h_{x}|h_{x^{\prime}}\rangle|\leq\eta$ for all pairs of distinct $\ell$-bit strings $x\neq x^{\prime}$. Then they argue that there exists a fixed $H$ which gives a quantum fingerprinting with the desired parameter by a non-constructive argument. We observe that the above construction can be directly used as quantum fingerprinting in the quantum random oracle model (QROM) [BDF+11]. This means that we have a quantum fingerprinting with output length $\log\ell+O(\log\eta^{-1})$ in the QROM, which is better than the one given in Lemma 2.13. We do not know how to achieve the same bound in the plain model (i.e., without relying on random oracles). The improvement from $O(\log(\ell)\log(\eta^{-1}))$ to $\log\ell+O(\log\eta^{-1})$ does not lead to any asymptotical improvement for output length of $\epsilon$-OWSGs in the parameter regime which we consider in this paper.
Lemma 2.15 ([Ras07, Lemma 4]).
For any Hermitian matrices $A$ and $B$, it holds that
$\max_{\Pi}\mathrm{Tr}(\Pi(A-B))=\|A-B\|_{tr}+\frac{\mathrm{Tr}(A-B)}{2}$
where the maximum is taken over all projections $\Pi$.
Corollary 2.16.
For any Hermitian matrices $A$ and $B$ and any projection $\Pi$, it holds that
$\left|\mathrm{Tr}\left(\Pi(A-B)\right)\right|\leq 2\|A-B\|_{tr}.$
Proof.
It immediately follows from Lemma 2.15 and $\mathrm{Tr}(M)\leq\|M\|_{1}=2\|M\|_{tr}$ for any Hermitian matrix $M$. ∎
Lemma 2.17.
Let {|$\{\ket{\psi_{k}}\}_{k\in\{0,1\}^{n}}$ be any set of pure states. Then
$F\left(\frac{1}{2^{n}}\sum_{k\in\{0,1\}^{n}}\ket{\psi_{k}}\bra{\psi_{k}},\frac{I^{\otimes m}}{2^{m}}\right)\leq\frac{2^{n}}{2^{m}}.$ (9)
Proof.
Let $\sum_{k\in\{0,1\}^{n}}\alpha_{k}\ket{\phi_{k}}\bra{\phi_{k}}$ | be a diagonalization of $\frac{1}{2^{n}}\sum_{k\in\{0,1\}^{n}}\ket{\psi_{k}}\bra{\psi_{k}}$ |. Then
$F\left(\frac{1}{2^{n}}\sum_{k\in\{0,1\}^{n}}\ket{\psi_{k}}\bra{\psi_{k}},\frac{I^{\otimes m}}{2^{m}}\right)=\left(\sum_{k}\sqrt{\frac{\alpha_{k}}{2^{m}}}\right)^{2}\leq(\sum_{k}\alpha_{k})\left(\sum_{k}\frac{1}{2^{m}}\right)=\frac{2^{n}}{2^{m}}.$ (10)
Here in the inequality we have used Cauchy–Schwarz inequality. ∎
3 Lower Bounds for ($\epsilon$-)OWSGs
In this section, we prove lower bounds for output length of ($\epsilon$-)OWSGs (Theorems 1.1, 1.2, 1.4 and 1.3).
3.1 Lower Bounds for OWSGs and $\lambda^{-\Theta(1)}$-OWSGs
@@
where Equation 20 follows from the correctness of $(\mathsf{KeyGen},\mathsf{StateGen},\mathsf{Ver})$ and Equation 11. Note that we have $\sqrt{1-x}\leq 1-x/2$ for all $0\leq x\leq 1$. Therefore we have
$\Pr[\mathcal{A}~{}\text{wins}]\geq 2^{-m-1}-{\mathsf{negl}}(\lambda).$ (21)
@@
The above proof can be easily extended to give a lower bound for $\lambda^{-\Theta(1)}$-OWSGs.
Theorem 3.3 (Restatement of Theorem 1.2).
For any constant $c\in\mathbb{N}$, there do not exist $\lambda^{-c}$-OWSGs of which output length is at most $c\log\lambda-2$.
Proof.
Suppose that the output length is $m<c^{\prime}\log\lambda-2$. By the proof of Theorem 3.1, there is a QPT adversary that wins with probability at least $2^{-m-1}-{\mathsf{negl}}(\lambda)\geq 2\lambda^{-c}-{\mathsf{negl}}(\lambda)\geq\lambda^{-c}$ for all sufficiently large $\lambda$. This means that it is not a $\lambda^{-c}$-OWSG. ∎
Remark 3.4.
As one can easily see from the proof, we can improve the lower bound from $c\log\lambda-2$ to $c\log\lambda-1-\epsilon$ for any constant $\epsilon>0$.
3.2 Lower Bounds for $\Theta(1)$-OWSGs and Weak OWSGs.
@@
Because of Lemmata 2.10 and 2.12, the time complexity of this algorithm is bounded by a polynomial in $1/\gamma,d,\lambda,T$ and $|\mathsf{Net}(d,\gamma)$. By the condition, $|\mathsf{Net}(d,\gamma)|\leq(C/\gamma)^{d^{2}}=(6C/\Delta)^{2^{2m}}={\mathrm{poly}}(\lambda)$ and this implies that $1/\Delta={\mathrm{poly}}(\lambda)$ and $d=2^{m}={\mathrm{poly}}(\lambda)$. This proves that $\mathcal{A}$ is a QPT algorithm.
@@
4 Upper Bounds for ($\epsilon$-)OWSGs
In this section, we prove upper bounds for output length of ($\epsilon$-)OWSGs (Theorems 1.5, 1.6 and 1.7).
4.1 Upper Bounds for OWSGs and $\lambda^{-\Theta(1)}$-OWSGs
@@
According to [BS20], if there exist OWFs, then there exist PRSGs with input length $\lambda$ and any output length $m$. From [CGG+23, Theorem 3.4], these PRSGs with $m=\omega(\log\lambda)$ are also OWSGs. ∎
We show that the result of [CGG+23] can be adapted to construct $1/{\mathrm{poly}}(\lambda)$-OWSGs with $O(\log\lambda)$-qubit outputs. We recall the following lemma from [CGG+23, Lemma 3.2] with a small modification.
Lemma 4.2.
Let $(\mathsf{KeyGen},\mathsf{StateGen})$ be an $\epsilon$-PRSG, where $\mathsf{StateGen}$ maps $n$-bit strings to $m$-qubit pure states. Let $h\in[0,1]$ and $\delta=2^{n}\cdot\left(1-h\right)^{2^{m}-1}+h$. Define a QPT algorithm $\mathsf{Ver}(k^{\prime},\ket{\phi_{k}})\to\top/\bot$ to be the projection of |$\ket{\phi_{k}}$ onto |$\ket{\phi_{k^{\prime}}}$ with the output $\top$ only when the projection is successful. Then $(\mathsf{KeyGen},\mathsf{StateGen},\mathsf{Ver})$ is an $(\epsilon+\delta)$-OWSG.
Proof sketch of Lemma 4.2.
Most parts of the proof are identical to the original one. When proving the security, we recall the following concentration inequality for Haar random states instead of [CGG+23, Lemma 3.1].
Lemma 4.3.
For any $h\in[0,1]$ and for any $m$-qubit quantum state $\phi_{0}$, it holds that for the Haar measure on $m$-qubit quantum states $\mu_{m}$,
$\Pr_{\ket{\psi}\leftarrow\mu_{m}}\left[|\braket{\psi}{\phi_{0}}|^{2}\geq h\right]=(1-h)^{2^{m}-1}.$
The proof of Lemma 4.3 can be found in the proof of [AK07, Lemma 3.6]. Using Lemma 4.3 instead of [CGG+23, Lemma 3.1] gives the choice $\delta=2^{n}\cdot(1-h)^{2^{m}-1}+h$, where $h$ has a similar role to $1/f(n)$ in the original proof. ∎
Theorem 4.4 (Restatement of Theorem 1.5).
For any constant $c\in\mathbb{N}$, there exist $\lambda^{-c}$-OWSGs with $(c+1)\log\lambda+O(1)$-qubit pure state outputs assuming the existence of OWFs.
Proof.
By [BS20], there exist PRSGs with input length $\lambda$ and any output length $m$ assuming the existence of OWFs. We apply Lemma 4.2 with $h=\lambda^{-c}/4$ and
$m\geq\log\left(\frac{\log_{e}\left(2^{\lambda+2}\lambda^{c}\right)}{h}+1\right)~{}\Longrightarrow~{}2^{n}\cdot(1-h)^{2^{m}-1}\leq\lambda^{-c}/4,$ (33)
which gives $\delta=\lambda^{-c}/2$. To prove the asymptotic bound of $m(\lambda)$, we choose a constant $K$ such that $2^{\lambda+2}\lambda^{c}\leq e^{\lambda}$ for all $\lambda\geq K$. In this case, the minimal $m$ satisfies
$m\geq\log\left(4\lambda^{c+1}+1\right)\geq(c+1)\log\lambda+3,$ (34)
thus we can choose $m(\lambda)=(c+1)\log\lambda+O(1)$. ∎
4.2 Upper Bound for $\Theta(1)$-OWSGs